OpenStudy (artemis17):

What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)?

OpenStudy (artemis17):

WILL GET MEDAL

OpenStudy (peachpi):

find the slope of the line first using the equation below. Pick one point to be \((x_1,y_1)\) and the other is \((x_2,y_2)\) and plug in the numbers \[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (artemis17):

okay so i did \[y=2-(-4)\div0-(-3)\]

OpenStudy (artemis17):

and got 6/3

OpenStudy (artemis17):

@peachpi

OpenStudy (peachpi):

yes, so that's 2.

OpenStudy (peachpi):

now plug into the point-slope formula \[y-y_1=m(x-x_1)\] m is the slope you just found you can use the same point you used above for (x1, y1)

OpenStudy (artemis17):

y - 8) = 2(x - 3) (y + 3) = -2(x + 4) (y - 2) = -2(x + 4) (y + 4) = 2(x + 3) these are the choices

OpenStudy (artemis17):

@peachpi

OpenStudy (peachpi):

y - (-4) = 2(x - (-3)) y + 4 = 2(x + 3)

OpenStudy (artemis17):

Okay, i think i got it now thanks!! @peachpi

OpenStudy (artemis17):

do you have time fore another one? it will be my last! @peachpi

OpenStudy (peachpi):

ok

OpenStudy (artemis17):

How do you find the rate of change for this question? @peachpi

OpenStudy (peachpi):

Use the slope equation we used before. Since they want the rate of change from 1 to 4, use those x and y values. At x = 1, y = 3, so that point is (1, 3). At x = 4, y = 0, so that point is (4, 0) \[\frac{ 0-3 }{ 4-1 }\]

OpenStudy (artemis17):

it says -1!

OpenStudy (artemis17):

so it would then be 0-3/4-(-1)?

OpenStudy (artemis17):

then after that it would then be... -3/5

OpenStudy (artemis17):

@peachpi

OpenStudy (peachpi):

it's -1

OpenStudy (peachpi):

the 1 in your denominator is positive

OpenStudy (artemis17):

oh my bad i was reading the question rather than what you wrote! @peachpi

OpenStudy (artemis17):

Okay thank you for help i understand noww!! It was correct!

OpenStudy (peachpi):

you're welcome